OpenStudy (anonymous):

Triangle ABC and DEF are congruent. Find x. triangle ABC: side AB measures x+6 side BC measures x side CA measures x+4 triangle DEF: side DE measures 3x-5

OpenStudy (anonymous):

Hun 12, 2013 - Side AB measures 3x – 5. Side BC measures 2x + 7. Side CD measures x + 12. Find the value of x. parallelogram A B C D Answer 12 8.5 5 2 . 2 points ... Triangle ABC is congruent to triangle XYZ. ... Triangle DEF is similar to triangle PQR. Side ... Side PQ measures 5x – 25 and side QR measures 6x + 4.

OpenStudy (anonymous):

Question 286031: The measures of the angles of a triangle are (2x + 10)◦, (3x)◦, (8x − 25)◦. ... Question 286562: triangle CAT has vertices c(4,9) A(8,-9) T(-6,5) ... The questions is: Find the length of the side opposite the 30 degree angle, x. .... It can be concluded that triangle ABC is congruent to triangle DEF, given BC = EF, ...

OpenStudy (anonymous):

Write an exact expression for the base and the height of the right triangle, ... Question 826917: The measures of the angles of the isosceles triangle are in the ratio 8:8:2 ... Question 825174: if the three sides of a triangle are 2, 5, and x units and ..... Question 824277: If triangle ABC is similar to triangle DEF, express AB in ...

OpenStudy (anonymous):

congruent sides measure (2x + 3) cm. ... Two sides of an equilateral triangle measure. (2y + 3) ... 4. AB. Find the total distance from A to B to C to D to E. 6. What additional ... ∆ABC is an isosceles triangle. ..... ∆DBC is an obtuse triangle by definition. ... Step 1 Find the value of x. BC. = CA. 8x + 3 = 7x + 8 x. = 5. Step 2 Find AB.

OpenStudy (mathmale):

Stacey: Looks like the others are doing OK in addressing your question. In the future, please consider graphing the triangles, as this would make it so much easier to visualize what's happening here.

OpenStudy (anonymous):

how would I submit it with a picture of it? @mathmale

OpenStudy (mathmale):



OpenStudy (anonymous):



OpenStudy (anonymous):

find x

OpenStudy (mathmale):

Nice drawing, Stacey. So much easier to read and understand than the purely verbal and numeric descriptions. Generally, when we're asked to "find x," we have to start with an equation. Here you have no equation, only expressions for the side lengths. Are you positive there was nothing else in the problem statement? For example, "the perimeter of DEF is 100 feet".

OpenStudy (anonymous):

Triangle ABC and DEF are congruent. Find x.

OpenStudy (anonymous):

that's the question copied and pasted

OpenStudy (mathmale):

Example: If the perimeter of DEF were 100 feet, then the perimeter of ABC would also be 100 feet, and the lengths of the sides of ABC would add up to 100 feet. Therefore, x+4 + x+6 + x =100 feet, and 3x+10=100, so x=30 feet.

OpenStudy (anonymous):

there is nothing in the problem saying anything else about or anything

OpenStudy (mathmale):

Side DE measures 3x-5 feet (not side EF, as you have it). But changing that won't affect the solution much, if at all, if we don't know the perimeter of one of the triangles.

OpenStudy (mathmale):

At this point I have no further insights to share. You might want to take up this discussion with others in your class, or, better yet, with your instructor. Would love to help more, but do not see anything I recognize that I haven't already shared with you.

OpenStudy (mathmale):

It may be more profitable for you to move on to other problems. perhaps you could come back to this one once you've completed the others.

OpenStudy (anonymous):

would this help?

OpenStudy (mathmale):

I appreciate your sharing this, but unfortunately cannot open this particular doc on my computer. What's in this doc?

OpenStudy (anonymous):

it was a screen shot of the problem exactly is there another way for me to do that so you can see it

OpenStudy (mathmale):

I'm running out of time right now. However, if you have a cell phone and are good at taking pictures with it, take one of the problem and forward the file to me at wrgnsc@rit.edu (as an attachment to e-mail).

OpenStudy (ranga):

Triangles ABC and DEF are congruent. Therefore, BC = EF. Equate the two and solve for x.

OpenStudy (ranga):

rather AB = DE as per OP as per drawing the above is correct.

OpenStudy (mathmale):

Ranga! Many thanks. Stacey: Wow. Why didn't I think of that? Solve x=3x-5.

OpenStudy (ranga):

np mathmale.

OpenStudy (mathmale):

StaceyMarie: x=3x-5 yields x = ? Notice, again, how much your drawing these triangles helped to clarify the situation we're dealing with.

OpenStudy (ranga):

Your original post says DE = 3x - 5. Your drawing says EF = 3x - 5. Which is correct?

OpenStudy (anonymous):

its EF=3x-5

OpenStudy (mathmale):

I'll let Stacey answer that. this teamwork has been very productive. Thank you again, ranga, for your help. Stacey: I have to get off the computer soon. But I hope to see y ou again soon on Open Study.

OpenStudy (anonymous):

oh okay @mathmale thank you so much

OpenStudy (mathmale):

You're most welcome! Good night!

OpenStudy (ranga):

Since ABC and DEF are congruent, BC = EF x = 3x - 5 subtract x from both sides: 0 = 2x - 5 add 5 to both sides: 5 = 2x divide both sides by 2: 2.5 = x

OpenStudy (ranga):

x = 2.5

OpenStudy (anonymous):

youre such a lifesaver! I have no idea what exactly had me stuck.

OpenStudy (ranga):

You are welcome. :)